# GUST: Quantifying Free-Form Geometric Uncertainty of Metamaterials Using Small Data

### v1

@@ 2.1 Past Work on Manufacturing Uncertainty Quantification
Figure 3: The forward diffusion process adds noise to data and the reverse process recovers the original data through iterative denoising. Here, $t=0,...,T$ is the timestep.
@@ 2.3 Transfer Learning for Generative Model
Early research on transfer learning in computer vision [17, 18] shows that early convolutional layers learn very general, low-level visual features—edges, corners, simple color or texture patterns—that are broadly useful across almost any image-classification task. Since these filters are not tied to specific object classes, we can safely freeze them and reuse their weights with little risk of overfitting. By contrast, deeper layers build up higher-level, task-specific patterns (object parts or whole-object prototypes) that vary more between datasets, so we typically fine-tune those to adapt the model to your new classes or domain. However, uncertainty in our problem could arise in both low- and high-level features. It is unclear at what level of features the difference exists between geometric uncertainties underlying the pretraining and fine-tuning data.
Transfer learning’s success in computer vision inspires its use in generative models, where limited data and overfitting pose similar challenges. To overcome the limitations of training generative models, transfer learning has been increasingly adopted to adapt pretrained models to new tasks. Training a generative model from scratch or fine-tuning a pretrained generative model on limited target domain data often leads to substantial performance decline due to overfitting and memorization [19]. This is mainly because the training data lacks diversity, failing to cover the full underlying data distribution. To address this, various studies have proposed methods for generative transfer learning. Specifically, transfer learning strategies for generative models can be grouped into two types: fine-tuning lightweight adapters [20, 21, 22] and fine-tuning with regularization [23, 19, 24, 25].
@@
Fine-tuning with regularization employs techniques like contrastive or adversarial losses to stabilize model adaptation and prevent overfitting when training on limited data. Zhang et al. [24] and Zhao et al. [25] focused on one-shot and few-shot generative tasks, respectively, using regularization techniques such as contrastive learning and adversarial losses to ensure stable and generalizable adaptation.
@@ 3.2 Self-Supervised Pretraining
$p_{\theta}(\mathbf{x}_{\text{fab},{t-1}}\mid\mathbf{x}_{\text{fab},t},\mathbf{x}_{\text{nom}})\\[-2.15277pt] =\mathcal{N}\bigl{(}\mathbf{x}_{\text{fab},{t-1}};\,\mu_{\theta}(\mathbf{x}_{\text{fab},t},t,\mathbf{x}_{\text{nom}}),\;\Sigma_{\theta}(\mathbf{x}_{\text{fab},t},t,\mathbf{x}_{\text{nom}})\bigr{)}$ (8)
@@
where $x_{t}=\sqrt{\bar{\alpha}_{t}}\,x_{0}+\sqrt{1-\bar{\alpha}_{t}}\,\epsilon$. In this work, we employ a conditional DDPM with concatenation embedding and SPADE layers [26] to learn the conditional distribution $p(\mathbf{x}_{\text{fab}}\mid\mathbf{x}_{\text{nom}})$. By concatenating the nominal geometry $\mathbf{x}_{\text{nom}}$ to intermediate feature maps at each diffusion step, the denoising network is guided to steer its reverse diffusion trajectory $p_{\theta}\bigl{(}\mathbf{x}_{\text{fab},\,t-1}\mid\mathbf{x}_{\text{fab},\,t},\mathbf{x}_{\text{nom}}\bigr{)}$ towards outputs that remain faithful to the condition. This process iteratively denoises random noise, conditioned on $\mathbf{x}_{\text{nom}}$, to capture the complex geometric variations of the fabricated geometry $\mathbf{x}_{\text{fab}}$.
@@ 3.4 Transfer Learning
This transfer learning process aims to bridge the gap between synthetic and real-world data. Knowing the level of features (low-level versus high-level) where the main geometric difference lies between the source and target domains informs the most suitable transfer learning strategy. For example, changing topology from the source to the target domains may require fine-tuning layers that capture high-level features. For the case where we have no prior knowledge regarding the nature of the target domain, the best approach is to try different transfer learning strategies, such as those introduced in Sec. 2.3. In our experiment, we explored several transfer learning strategies, among which no obvious distinctions were observed in their performance. The impact of transfer learning strategies on the generative performance will be studied in future work.
@@ 3.5 Uncertainty Quantification
To quantify the geometric uncertainty of fabricated metamaterials, we use the resulting generative model to generate Monte Carlo samples of as-fabricated geometries given the target nominal design as the input condition, i.e., $\mathbf{x}_{\text{fab}}^{(i)}\sim p_{\theta}(\mathbf{x}_{\text{fab}}\mid\mathbf{x}_{\text{nom}})$.
@@ 4.1 Data Preparation
We used unit cell geometries selected from the 2D orthotropic metamaterial microstructure dataset [27] as nominal designs, and created the pretraining dataset $\mathcal{D}_{\text{pretrain}}$ by applying FFD (Eq. (6)) to the nominal geometries. Specifically, each nominal design was deformed into 20 distinct variants, resulting in a total of 60,000 as-fabricated geometries derived from 3,000 unique nominal designs. We added randomness to the as-fabricated geometries by adding Gaussian noise with a standard deviation $\sigma=6$ to the FFD control points. All geometries were represented as binary pixelated images with a resolution of $64\times 64$.
@@ 4.3 Pretraining and Transfer Learning
We then fine-tuned the pretrained model on $\mathcal{D}_{\text{finetune}}$. We experimented with multiple configurations in this study by transferring weights and biases from different layers of the pretrained model (i.e., freezing these layers’s parameters during fine-tuning). We fine-tuned each model for 38,400 iterations, the learn rate will decay by 10% every 960 iterations with the same batch size as in pretraining. Two distinct approaches were investigated as follows.
@@ 4.4 Baselines
The dilation-erosion approach [6] (Fig. 2, top left) approximates manufacturing uncertainty by complementary morphological operations that respectively expand and contract shape boundaries using a predefined structuring element. We treat the sizes of the dilation and erosion operators as tunable parameters and choose them by using maximum likelihood estimation. Given a set of nominal designs $X=\{\mathbf{x}_{\text{nom},i}\}_{i=1}^{N}$ and a reference collection of real-world fabricated image $\{\mathbf{x}_{\text{fab},j}\}_{j=1}^{M}$, we define a nonparametric Gaussian-kernel density estimator:
@@
where $h$ is the kernel bandwidth. For a positive integer dilation scale $\alpha$, we apply the operator $\mathcal{F}_{\alpha}$, which expands shape boundaries in an image by applying a morphological dilation with a square kernel of size $\alpha$, to each nominal design. The log-likelihood can be computed as
@@
We then obtain $\hat{\alpha}=\arg\min_{\alpha}\,\mathcal{-L}_{d}(\alpha)$ by minimizing the negative log-likelihood via differential evolution [30]. Similarly, the log-likelihood corresponding to the erosion operator $\mathcal{F}_{\beta}$, which contracts shape boundaries using a morphological erosion with a square kernel of size $\beta$, can be expressed as
@@
yields the optimal erosion scale $\hat{\beta}=\arg\min_{\beta}\,\mathcal{-L}_{e}(\beta)$. This procedure produces the dilation and erosion sizes that make the processed geometries closest to real-world samples, serving as a principled baseline for comparison.
@@
This approach models geometric variability by adding a GRF to the signed distance function (SDF) of the nominal design [4]. Let $\{\mathbf{x}_{i}\}_{i=1}^{N}$ be points on a uniform grid within $[0,1]^{2}$, and define the precision matrix
[…]
@@
We perform the eigen-decomposition $C\phi_{k}=\lambda_{k}\phi_{k}$, sort the eigenvalues $\{\lambda_{k}\}$ in descending order, and truncate the decomposition to retain only the top $M$ modes.
A single GRF realization is generated via the Karhunen–Loève (KL) expansion,
@@
After normalizing $g$ to unit standard deviation, we scale by amplitude $\alpha$ and add it to the original SDF $s(\mathbf{x}_{i})$:
$s_{\mathrm{noisy}}(\mathbf{x}_{i})=s(\mathbf{x}_{i})+\alpha\,\frac{g(\mathbf{x}_{i})}{\sigma(g)},$
where $\sigma(g)$ is the standard deviation of $g$. While the GRF approach offers a significantly higher degree of freedom in modeling geometric variability compared to dilation-erosion, it still relies on simplifying assumptions inherent in its Gaussian nature and KL expansion.
@@ 4.5 Quantitative Metrics
To mitigate the computational expense of computing the density and coverage metrics in the original high-dimensional pixel space, we employed t-distributed stochastic neighbor embedding (t-SNE) with a perplexity of 10 for dimensionality reduction. This projection into a three-dimensional embedding space preserves local neighbor relationships, a crucial requirement for the accurate calculation of the density and coverage metrics.
@@
Figure 9: KDEs of homogenized elastic properties ($C_{11}$, $C_{12}$, $C_{22}$, $C_{33}$) for the synthetic experiment. The corresponding normal designs are shown at the top left corners. GUST produces a property distribution either closely matching the ground truth or situated between the distributions associated with the source domain and the ground truth.
@@ 5.1 Synthetic Experiment
In addition to generated geometries, we analyzed the elastic properties of geometries generated by GUST and baseline approaches. Figure 8 presents the distributions of homogenized elastic tensor components ($C_{11}$, $C_{12}$, $C_{22}$, $C_{33}$) corresponding to two nominal designs, where the “ground truth” is based on effective properties computed from the geometries in $\mathcal{D}_{\text{finetune, synth}}$. While we picked these representative results, more results are available in Appendix A. Notably, the dilation-erosion baseline can only lead to point estimates instead of a full distribution. We employed MLE to determine both the dilation and erosion scales, as introduced in Sec. 4.4. However, the MLE result for the erosion scale was zero, indicating that the fabricated geometries predominantly exhibit a dilated pattern. Despite optimizing the dilation process using MLE, it still fails to align with even the mean of the ground truth distribution for most cases. This is likely due to MLE’s reliance solely on geometric properties rather than the underlying material properties. The GRF approach typically exhibits a markedly different distribution compared to the ground truth, with some exceptions shown in Appendix A. Given the inherently random nature of GRF, its occasional close alignment with ground truth is expected. Compared to GUST, the distributions corresponding to DT are generally farther from the ground truth. DT’s performance can be extremely unstable (e.g., Design 2 in Fig. 8), possibly due to training instability with limited data. Overall, the GUST shows a relatively stable performance, evident by its consistently close alignment with ground truth across most of the components and designs. These qualitative observations will be further supported by quantitative results via density, coverage, and Wasserstein distance assessment.
@@
We conducted a quantitative analysis of the geometric quality of generated samples through the density and coverage metrics (Sec. 4.5). Figure 10 shows the distribution of these metrics evaluated on 20 nominal designs, each with 3,000 samples of as-fabricated geometries, ensuring a thorough assessment of the models’ performance across diverse scenarios. The density metric quantifies how closely the generated samples cluster around the real data points, while the coverage metric measures how broadly the generated samples span the real data distribution. GUST achieved substantially higher mean density with lower variance than GRF and DT. The p-values in Table 1 indicate a significant difference in the means of the density metric between GUST and the other two baselines. Meanwhile, GUST’s coverage metric differs only marginally from GRF (mean values differ by 0.07). Based on Fig 10, both GUST and GRF achieve a moderate coverage, meaning that neither of them matches the ground truth well in terms of the distribution of geometries in the reduced space. Considering GUST’s generally higher density metrics, we infer that while its high-density regions may overlap with the ground truth, it tends to produce a narrower distribution that lacks coverage. In contrast, GRF receives median scores on both metrics, indicating a possible shift in its distribution’s mode compared to the ground truth. A similar coverage to GUST is possibly attributed to GRF’s broader distribution. This observation is consistent with the property distributions in Fig 8.
Figure 10: Density and coverage (higher is better) metrics for geometries generated by different methods in the synthetic experiment, calculated across 20 nominal designs (each with 3,000 samples of as-fabricated geometries). Density measures sample clustering around real data, while coverage assesses the span of generated samples.
Figure 11: Wasserstein distances (lower is better) between property distributions from different approaches and the ground truth distribution of elastic tensor components ($C_{11}$, $C_{12}$, $C_{22}$, $C_{33}$), calculated across 20 nominal designs (each with 3,000 samples of as-fabricated geometries) in the synthetic experiment. GUST consistently shows lower average Wasserstein distances and significantly lower variance.
@@
We use the Wasserstein distance to measure the alignment between the property distributions of geometries generated by different approaches and the ground truth. Wasserstein distances were calculated using the same generated geometries as those used to calculate the density and coverage metrics (i.e., 20 nominal designs, each with 3,000 generated as-fabricated geometries). Figure 11 shows the distributions of calculated Wasserstein distances for the elastic tensor components ($C_{11}$, $C_{12}$, $C_{22}$, $C_{33}$). GUST consistently achieves lower mean values with significantly lower variances. The p-values in Table 1 further demonstrate the significant difference in the mean Wasserstein distances between GUST and the two baselines.

### v2

@@ 2.1 Past Work on Manufacturing Uncertainty Quantification
Figure 3: The forward diffusion process adds noise to data and the reverse process recovers the original data through iterative denoising. Here, t=0,,Tt=0,...,T is the timestep.
@@ 2.3 Transfer Learning for Generative Model
Early research on transfer learning in computer vision [17, 18] shows that early convolutional layers learn very general, low-level visual features—edges, corners, simple color or texture patterns—that are broadly useful across almost any image-classification task. Since these filters are not tied to specific object classes, we can safely freeze them and reuse their weights with little risk of overfitting. By contrast, deeper layers build up higher-level, task-specific patterns (object parts or whole-object prototypes) that vary more between datasets, so we typically fine-tune those to adapt the model to new classes or domain. However, uncertainty in our problem could arise in both low- and high-level features. It is unclear at what level of features the difference exists between geometric uncertainties underlying the pretraining and fine-tuning data.
Transfer learning’s success in computer vision inspires its use in generative models, where limited data and overfitting pose similar challenges. To overcome the limitations of training generative models, transfer learning has been increasingly adopted to adapt pretrained models to new tasks. Training a generative model from scratch or fine-tuning a pretrained generative model on limited target domain data often leads to substantial performance decline due to overfitting and memorization [19]. This is mainly because the training data lacks diversity, failing to cover the full underlying data distribution. To address this, various studies have proposed methods for generative transfer learning. Specifically, transfer learning strategies for generative models can be grouped into two main types: fine-tuning lightweight adapters [20, 21, 22] and fine-tuning with regularization [23, 19, 24, 25].
@@
Fine-tuning with regularization employs techniques like contrastive or adversarial losses to stabilize model adaptation and prevent overfitting when training on limited data. For example, Zhang et al. [24] and Zhao et al. [25] focused on one-shot and few-shot generative tasks, respectively, using regularization techniques such as contrastive learning and adversarial losses to ensure stable and generalizable adaptation. Regularizers such as L2-Starting-Point (L2-SP) [26] were proposed to avoid large deviations from the initial pretrained weights and mitigate catastrophic forgetting.
Some studies specifically focus on conditional generative models or diffusion models. In particular, Cheng et al. [27] showed a theoretical study on how the representation of conditions learned from source tasks can improve the sample efficiency of target tasks in transfer learning. Ouyang et al. [28] proposed modeling the score function on the target domain as the score function on the source domain with additional guidance. The guidance network learns the density ratio of the target and source domain data distributions via a binary classifier. Huijben et al. [29] proposed unconditional pre-training to learn general representations, followed by conditional fine-tuning. The conditions (labels) are only incorporated in the fine-tuning stage by adding the trainable label embeddings to time step embeddings.
@@ 3.2 Self-Supervised Pretraining
$p_{\theta}(\mathbf{x}_{\text{fab},{t-1}}\mid\mathbf{x}_{\text{fab},t},\mathbf{x}_{\text{nom}})=\mathcal{N}\bigl{(}\mathbf{x}_{\text{fab},{t-1}};\,\mu_{\theta}(\mathbf{x}_{\text{fab},t},t,\mathbf{x}_{\text{nom}}),\;\Sigma_{\theta}(\mathbf{x}_{\text{fab},t},t,\mathbf{x}_{\text{nom}})\bigr{)}$ (8)
@@
where $x_{t}=\sqrt{\bar{\alpha}_{t}}\,x_{0}+\sqrt{1-\bar{\alpha}_{t}}\,\epsilon$. In this work, we employ a conditional DDPM with concatenation embedding and SPADE layers [30] to learn the conditional distribution $p(\mathbf{x}_{\text{fab}}\mid\mathbf{x}_{\text{nom}})$. By concatenating the nominal geometry $\mathbf{x}_{\text{nom}}$ to intermediate feature maps at each diffusion step, the denoising network is guided to steer its reverse diffusion trajectory $p_{\theta}\bigl{(}\mathbf{x}_{\text{fab},\,t-1}\mid\mathbf{x}_{\text{fab},\,t},\mathbf{x}_{\text{nom}}\bigr{)}$ towards outputs that remain faithful to the condition. This process iteratively denoises random noise, conditioned on $\mathbf{x}_{\text{nom}}$, to capture the complex geometric variations of the fabricated geometry $\mathbf{x}_{\text{fab}}$. The pseudocode of self-supervised pretraining is shown in Algorithm 1 of Appendix C.
@@ 3.4 Transfer Learning
This transfer learning process aims to bridge the gap between synthetic and real-world data. In this stage, we fine-tune the pretrained conditional DDPM on the real-world dataset. The pseudocode of fine-tuning is shown in Algorithm 2 of Appendix C. Knowing the level of features (low-level versus high-level) where the main geometric difference lies between the source and target domains informs the most suitable transfer learning strategy. For example, changing the topology from the source to the target domains may require fine-tuning layers that capture high-level features. For the case where we have no prior knowledge regarding the nature of the target domain, the best approach is to try different transfer learning strategies, such as those introduced in Sec. 2.3. In our experiment, we explored several transfer learning strategies, among which no obvious distinctions were observed in their performance. Descriptions of the transfer learning strategies and their quantitative results are shown in Appendix B. Future work will explore additional transfer learning strategies.
@@ 3.5 Uncertainty Quantification
To quantify the geometric uncertainty of fabricated metamaterials, we use the resulting generative model to generate Monte Carlo samples of as-fabricated geometries given the target nominal design as the input condition, i.e., $\mathbf{x}_{\text{fab}}^{(i)}\sim p_{\theta}(\mathbf{x}_{\text{fab}}\mid\mathbf{x}_{\text{nom}})$. Algorithm 3 in Appendix C shows the pseudocode for sampling the fine-tuned DDPM.
@@ 4.1 Data Preparation
We used unit cell geometries selected from the 2D orthotropic metamaterial microstructure dataset [31] as nominal designs, and created the pretraining dataset $\mathcal{D}_{\text{pretrain}}$ by applying FFD (Eq. (6)) to the nominal geometries. Specifically, each nominal design was deformed into 20 distinct variants, resulting in a total of 60,000 as-fabricated geometries derived from 3,000 unique nominal designs. While this small number of as-fabricated geometries can be insufficient for independently learning each conditional distribution $p(\mathbf{x}_{\text{fab}}|\mathbf{x}_{\text{nom}})$, the conditional DDPM learns from as-fabricated geometries across 3,000 nominal designs. This allows it to capture the covariance between the as-fabricated geometries of different nominal designs, therefore enabling robust learning through generalization, even with limited data per nominal design.
We added randomness to the as-fabricated geometries by adding Gaussian noise with a standard deviation $\sigma=6$ to the FFD control points. All geometries were represented as binary pixelated images with a resolution of $64\times 64$.
@@ 4.3 Pretraining and Transfer Learning
We then fine-tuned the pretrained model on $\mathcal{D}_{\text{finetune}}$. We experimented with multiple configurations in this study by transferring weights and biases from different layers of the pretrained model (i.e., freezing these layers’ parameters during fine-tuning). We fine-tuned each model for 38,400 iterations, the learning rate will decay by 10% every 960 iterations with the same batch size as in pretraining. Two distinct approaches were investigated as follows.
@@ 4.4 Baselines
The dilation-erosion approach [6] (Fig. 2, top left) approximates manufacturing uncertainty by complementary morphological operations that respectively expand and contract shape boundaries using a predefined structuring element. The expansion (dilation) and contraction (erosion) operators $\mathcal{F}_{\alpha}$ and $\mathcal{F}_{\beta}$ are realized by convolving $\mathbf{x}_{\text{nom}}$ with a square kernel of sizes $\alpha$ and $\beta$, respectively. During the convolution, we compute the maximal (minimal) pixel value overlapped by the kernel and replace the image pixel in the anchor point position with that maximal (minimal) value for dilation (erosion). The parameters $\alpha$ and $\beta$ control how much a geometry dilates or erodes, respectively. In this work, we implement $\mathcal{F}_{\alpha}$ and $\mathcal{F}_{\beta}$ using OpenCV’s dilate and erode methods. One can also use the approach detailed in Ref. [6] to implement the two operators. To ensure a fair comparison, we treat the $\alpha$ and $\beta$ in the dilation and erosion operators as tunable parameters and choose them by using maximum likelihood estimation (MLE), so that the operators’ effects accurately reflect the characteristics inherent in the data. Specifically, given a set of nominal designs $X=\{\mathbf{x}_{\text{nom},i}\}_{i=1}^{N}$ and a reference collection of real-world fabricated image $\{\mathbf{x}_{\text{fab},j}\}_{j=1}^{M}$, we define a Gaussian kernel density estimator:
@@
where $h$ is the kernel bandwidth. Given a positive integer dilation scale $\alpha$, we apply $\mathcal{F}_{\alpha}$ to each nominal design $\mathbf{x}_{\text{nom},i}$. The log-likelihood can be computed as
@@
We then obtain the maximum likelihood estimate $\hat{\alpha}=\operatorname*{arg\,min}_{\alpha\in[0,100]}\,\mathcal{-L}_{d}(\alpha)$ by minimizing the negative log-likelihood via differential evolution [34]. Similarly, the log-likelihood corresponding to the erosion can be expressed as
@@
which yields the maximum likelihood estimate $\hat{\beta}=\operatorname*{arg\,min}_{\beta\in[0,100]}\,\mathcal{-L}_{e}(\beta)$. This procedure produces the dilation and erosion kernel sizes that make the processed geometries closest to real-world distributions, serving as a principled baseline for comparison.
@@
This approach models geometric variability by adding a GRF to the signed distance function (SDF) of the nominal design [4]. Let $\{\mathbf{x}_{i}\}_{i=1}^{N}$ be points on a uniform grid within $0,1]^{2}$, and define the precision matrix
$L^{-1}=\mathrm{diag}\!\bigl{(}1/\ell_{1}^{2},\;1/\ell_{2}^{2}\bigr{)}.$
@@
We perform the eigen-decomposition $C\phi_{k}=\lambda_{k}\phi_{k}$, sort the eigenvalues $\{\lambda_{k}\}$ in descending order, and truncate the decomposition to retain only the top $M$ modes. A single GRF realization is generated via the Karhunen–Loève (KL) expansion,
@@
While there are non-Gaussian models such as the field of Gaussian scale mixtures (FoGSM) [35], testing on these models is beyond the scope of this work. In particular, we can place model assumptions along a spectrum. The dilation-erosion and GRF assumptions are at one end of the spectrum, while the completely “free-form” geometric variation learned by a deep generative model is at the other end. The non-Gaussian model that can learn from data is in the middle of the spectrum—it has higher flexibility than the strong dilation-erosion and GRF assumptions, yet is not completely “free-form” and therefore still risks making incorrect assumptions compared to a deep generative model. On the other hand, we need the model prediction to generalize to different nominal designs, i.e., given an unseen nominal design, the model needs to predict the corresponding geometric uncertainty without the data of as-fabricated geometries. However, a non-Gaussian model typically does not allow such generalizability without more advanced methodology development.
@@ 4.5 Quantitative Metrics
To mitigate the computational expense of computing the density and coverage metrics in the original high-dimensional pixel space, we employed t-distributed stochastic neighbor embedding (t-SNE) with a perplexity of 10. This perplexity was chosen as it led to the most reasonable representation of the local and global structures of real and generated data. This projection into a three-dimensional embedding space preserves local neighbor relationships, a crucial requirement for the accurate calculation of the density and coverage metrics.
@@ 5.1 Synthetic Experiment
Figure 9: KDEs of homogenized elastic properties for the synthetic experiment showing the discrepancy between the target domain, the source domain, and the distribution learned by GUST. GUST produces a property distribution either closely matching the ground truth or situated between the distributions associated with the source domain and the ground truth.
In addition to generated geometries, we analyzed the elastic properties of geometries generated by GUST and baseline approaches. Figure 8 presents the distributions of homogenized elastic tensor components ($C_{11}$, $C_{12}$, $C_{22}$, $C_{33}$) corresponding to two nominal designs, where the “ground truth” is based on effective properties computed from the geometries in $\mathcal{D}_{\text{finetune, synth}}$. While we picked these representative results, more results are available in Appendix A. Notably, the dilation-erosion baseline can only lead to point estimates instead of a full distribution. We employed MLE to determine both the dilation and erosion scales, as introduced in Sec. 4.4. The resulting dilation and erosion scales are 2 and 1, respectively. Note that a scale of 1 means no change to the nominal geometry. Despite optimizing the dilation scale using MLE, it consistently overestimates the actual elastic components (Figs. 8 and 16), possibly due to the ignorance of topological changes (i.e., random holes). The GRF approach typically exhibits a markedly different distribution compared to the ground truth, with some exceptions shown in Appendix A. Given the inherently random nature of GRF, its occasional close alignment with ground truth is expected. Compared to GUST, the distributions corresponding to DT are generally farther from the ground truth. DT’s performance can be extremely unstable (e.g., Design 2 in Fig. 8), possibly due to training instability with limited data. Overall, the GUST shows a relatively stable performance, evident by its consistently close alignment with ground truth across most of the components and designs. These qualitative observations will be further supported by quantitative results via density, coverage, and Wasserstein distance assessment.
@@
Figure 10: Density and coverage (higher is better) metrics for geometries generated by different methods in the synthetic experiment, calculated across 30 nominal designs (each with 3,000 samples of as-fabricated geometries). Density measures sample clustering around real data, while coverage assesses the span of generated samples.
We conducted a quantitative analysis of the geometric quality of generated samples through the density and coverage metrics (Sec. 4.5). Figure 10 shows the distribution of these metrics evaluated on 30 nominal designs, each with 3,000 samples of as-fabricated geometries, ensuring a thorough assessment of the models’ performance across diverse scenarios. The density metric quantifies how closely the generated samples cluster around the real data points, while the coverage metric measures how broadly the generated samples span the real data distribution. GUST achieved substantially higher mean density with lower variance than GRF and DT. The p-values in Table 1 indicate a significant difference in the means of the density metric between GUST and the other two baselines. Meanwhile, GUST’s coverage metric differs only marginally from GRF (mean values differ by 0.07). Based on Fig 10, both GUST and GRF achieve a moderate coverage, meaning that neither of them matches the ground truth well in terms of the distribution of geometries in the reduced space. Considering GUST’s generally higher density metrics, we infer that while its high-density regions may overlap with the ground truth, it tends to produce a narrower distribution that lacks coverage. In contrast, GRF receives median scores on both metrics, indicating a possible shift in its distribution’s mode compared to the ground truth. A similar coverage to GUST is possibly attributed to GRF’s broader distribution. This observation is consistent with the property distributions in Fig 8.
Figure 11: Wasserstein distances (lower is better) between property distributions from different approaches and the ground truth distribution of elastic tensor components ($C_{11}$, $C_{12}$, $C_{22}$, $C_{33}$), calculated across 30 nominal designs (each with 3,000 samples of as-fabricated geometries) in the synthetic experiment. GUST consistently shows lower average Wasserstein distances and significantly lower variance.
@@
We use the Wasserstein distance to measure the alignment between the property distributions of geometries generated by different approaches and the ground truth. Wasserstein distances were calculated using the same generated geometries as those used to calculate the density and coverage metrics (i.e., 30 nominal designs, each with 3,000 generated as-fabricated geometries). Figure 11 shows the distributions of calculated Wasserstein distances for the elastic tensor components ($C_{11}$, $C_{12}$, $C_{22}$, $C_{33}$). GUST consistently achieves lower mean values with significantly lower variances. The p-values in Table 1 further demonstrate the significant difference in the mean Wasserstein distances between GUST and the two baselines.